a car uses 6 gallons of gasolonie on a trip of 126miles. at that rate, how much gasoline will a trip of 294 miles require?

Answer by checkley71(8403)

126/6=294/x cross multiply
126x=6*294
126x=1764
x=1764/126
x=14 gallons of gas will be required to drive 294 miles.